OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y

OpenStudy (anonymous):

x – 2 = – 2y Solve for \(x\). \[x = -2y + 2\]

OpenStudy (anonymous):

Now input that value into the first equation.

OpenStudy (anonymous):

\[4-2y + 2 + 2y = 11\] Move all the like terms to one side.

OpenStudy (anonymous):

@zobobozo Can you do that?

OpenStudy (anonymous):

Wait, sorry forgot the round brackets... \[4(-2y + 2) + 2y = 11\]

OpenStudy (anonymous):

Is that what you got? Odd I got: \[x = 3,~y = \dfrac{-1}{2}\]

OpenStudy (anonymous):

\[4(-2y + 2) + 2y = 11\] Simplifies to: \[y = \dfrac{-1}{2}\]

OpenStudy (anonymous):

Then: \[x – 2 = – 2y\] Input -1/2: \[x – 2 = – 2\dfrac{-1}{2}\] Simplifies to: \[x = 3\]

OpenStudy (anonymous):

no its 3 and -1/2

OpenStudy (anonymous):

Uh, I said that? \[x = 3\] \[y = \dfrac{-1}{2}\] If you wanted to write that in a pair it would look like: \[3, \dfrac{-1}{2}\]